sin ​​​​×+cos*=1

sin​​​​×+cos*=1

Grade:12th pass

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:Hello student, please find answer to your question
Please rewrite the complete question.
Properly write the all the functions.
Complete the equation.
Also, define the interval in which you want the results.

sin(x) + cos(x) = 1
sin(a+b) = sin(a)cos(b) + cos(a)sin(b)
sin(x+\frac{\pi }{4}) = sin(x)cos(\frac{\pi }{4}) + cos(x)sin(\frac{\pi }{4})
sin(x+\frac{\pi }{4}) = \frac{sin(x) + cos(x)}{\sqrt{2}}
sin(x) + cos(x)=\sqrt{2}sin(x+\frac{\pi }{4})
\sqrt{2}sin(x+\frac{\pi }{4}) = 1
sin(x+\frac{\pi }{4}) = \frac{1}{\sqrt{2}}
x+\frac{\pi }{4} = 2n\pi + \frac{\pi }{4}
x = 2n\pi
x = 0, 2\pi